Write an equation of the line containing the given point and parellel to the given line. Express your answer in the form y = mx + b
(5,8); x + 2y = 5
the equation of the line is y= ?

Answer by ankor@dixie-net.com(22740)

Write an equation of the line containing the given point and parallel to the given line.
Express your answer in the form y = mx + b
(5,8); x + 2y = 5
Find the slope (parallel lines have the same slope)
x + 2y = 5
2y = -x + 5
y = -.5x + 2.5
slope (m) = -.5
the equation of the line is y= ?
:
y - 8 = -.5(x - 5)
y - 8 = -.5x + 2.5
y = -.5x + 2.5 + 8
y = -.5x + 10.5 is the equation of the line